Consider two neighboring island countries called Felicidad and Contente. They each have 4 million labor hours available per week that they can use to produce corn, jeans, or a combination of both. The following table shows the amount of corn or jeans that can be produced using 1 hour of labor.


1
Expert's answer
2021-12-27T08:59:02-0500

Solution:


Dolorium's opportunity cost of producing 1 pair of jeans is ("\frac{16bushles}{4pairs}" = 4 bushels of rye), and


Arcadia's opportunity cost of producing 1 pair of jeans is ("\frac{12bushles}{6pairs}" = 2 bushels of rye).


Similarly, Dolorium's opportunity cost of producing 1 bushel of rye is ("\frac{4 pairs}{16bushles}" = 0.25 pairs of

jeans),

and Arcadia's opportunity cost of producing one bushel of rye is ("\frac{6pairs}{12bushles}" = 0.5 pairs of jeans).


Therefore, Arcadia has a comparative advantage in the production of jeans, and Dolorium has a comparative advantage in the production of rye.

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces jeans will produce (6 pairs x 4 million hours = 24 million pairs per week), and the country that produces rye will produce (16 pairs x 4 million hours = 64 million bushels per week).

This is so because, after specialization, each country will divert its total available labor hours to the production of the specialized good which it needs to trade with the other country.